# Polymatroid Schubert varieties

### v1

@@ 3. Combinatorial flats
Our results on polymatroid Schubert varieties depend only on Sections 3.1 and 3.2, so geometrically-inclined readers may go directly to Section 4 after reading these.
@@
A combinatorial flat of $(P,\mathbf{n})$ is a multiset $\mathbf{s}\in\mathbb{Z}_{\geq 0}^{E}$ that is represented by a flat of $\widetilde{P}$.
@@
Suppose that $P$ is realized by a subspace arrangement $V_{1},\ldots,V_{N}\subset V$. Rephrasing 2.10 in the language of arrangements, if $\mathbf{n}$ is a cage for $P$, then $\widetilde{P}$ is realized by any arrangement $\{H_{ij}:1\leq i\leq N,1\leq j\leq n_{i}\}$ in which $H_{i1},\ldots,H_{in_{i}}$ are general hyperplanes containing $V_{i}$. The lattice of flats $\mathcal{L}_{\widetilde{P}}$ has an action by a product of symmetric groups $\mathfrak{S}$, and the quotient $\mathcal{L}_{\widetilde{P}}/\mathfrak{S}$ is isomorphic to $\mathcal{L}_{P}$ (see also Section 3.1).
@@
Let $i\in E$. The combinatorial flats of $(P\setminus i,\pi_{E\setminus i}(\mathbf{n}))$ are of the form $\pi_{E\setminus i}(\mathbf{s})$ with $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$.
Proof.
Apply 2.11 and 2.12. ∎
Lemma 3.10.
@@
The “if” direction is 3.9. For the “only if”, suppose that $\pi_{E\setminus i}(\mathbf{s})$ is a combinatorial flat of $(P\setminus i,\pi_{E\setminus i}(\mathbf{n}))$. We will show that $\mathbf{s}$ must be a flat by proving $\mathbf{s}=\overline{\mathbf{s}}$. By 3.2 and the hypothesis that $\overline{s}_{i}<n_{i}$, it is always the case that $s_{i}=\overline{s}_{i}$. Hence, it suffices to show that $s_{j}=\overline{s}_{j}$ for $j\neq i$.
By 3.9, there is a combinatorial flat $\mathbf{s}^{\prime}$ of $(P,\mathbf{n})$ with $\pi_{E\setminus i}(\mathbf{s}^{\prime})=\pi_{E\setminus i}(\mathbf{s})$. If $s^{\prime}_{i}=n_{i}$, then $\mathbf{s}<\mathbf{s}^{\prime}$, so $\mathbf{s}\leq\overline{\mathbf{s}}<\mathbf{s}^{\prime}$. Entrywise, this implies $s_{j}\leq\overline{s}_{j}\leq s_{j}^{\prime}=s_{j}$, which gives the desired equality.
@@
By the matroid case and 2.14, a set $S$ is a flat of $\widetilde{T_{F}P}$ if and only if $S$ is a flat of $\widetilde{P}$ that either contains $\cup_{i\in F}E_{i}$ or satisfies $\overline{S\cup i}\not\supset\cup_{i\in F}E_{i}$ for all $i\in\widetilde{E}\setminus S$. In the former case, $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)-1$; in the latter $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)$. Taking $\mathbf{s}$ to be the multiset represented by $S$ completes the proof. ∎
In the remainder of Section 3, we work up to proving 1.1, then provide axioms for combinatorial flats. These results are not required to understand polymatroid Schubert varieties; readers interested solely in geometry may proceed to Section 4 forthwith.
@@
$\leq\operatorname{rk}_{\widetilde{P}}(S)+\operatorname{rk}_{\widetilde{P}}(S^{\prime})=\operatorname{rk}_{P}(\mathbf{s})+\operatorname{rk}_{P}(\mathbf{s}^{\prime}).$
@@
$\mathcal{L}_{P,\mathbf{n}}$ is a top-heavy semimodular lattice, with join and meet given by $\vee$ and $\wedge$, respectively.
@@
where $S$ is any flat representing $\mathbf{s}$. By [BHM+20b, Theorem 1.1], there is an injective linear $L:H^{k}(\widetilde{P})\to H^{d-k}(\widetilde{P})$ that commutes with the action of $\mathfrak{S}$. Consequently, the restriction of $L$ to $H^{k}(\widetilde{P})^{\mathfrak{S}}$ gives a linear injection
@@
and taking dimensions yields top-heaviness. ∎
Generalizations of Lemmas 3.12 and 3.13 for matricubes appear in [AG24, Theorem 3.4] (see 3.8).
@@
By 3.14, each $i\in E$ corresponds to a maximal join-irreducible $x_{i}$ of $\mathcal{L}_{P,\mathbf{n}}$. The rank function of $P$ is recovered by $\operatorname{rk}_{P}(S)=\operatorname{rk}_{P,\mathbf{n}}(\vee_{i\in S}x_{i})$. ∎
@@
is an isomorphism. Multisets in the image of $\psi$ are combinatorial flats of $R_{i}(P,\mathbf{n})$ by Lemmas 2.16 and 2.11.
For surjectivity, suppose that $\mathbf{s}\in\mathcal{L}_{R_{i}(P,\mathbf{n})}$ is represented by a flat $S$ of $\widetilde{R_{i}(P,\mathbf{n})}$. We will show there is $\mathbf{s}^{\prime}\in\mathcal{L}_{P,\mathbf{n}}$ with $\psi(\mathbf{s}^{\prime})=\mathbf{s}$. By Lemmas 2.16 and 2.11, there is $e\in E_{i}\setminus S$ such that $S\cup e$ or $S$ is a flat of $\widetilde{P}$. If $S$ is a flat of $\widetilde{P}$, then $s_{i}<n_{i}$, so we may take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, suppose $S\cup e$ is a flat of $\widetilde{P}$. If $s_{i}+1<n_{i}$, then by 3.2, $\mathbf{s}$ is a combinatorial flat of $\widetilde{P}$, so we may again take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, if $S\cup e$ is a flat of $\widetilde{P}$ and $s_{i}+1=n_{i}$, then take $\mathbf{s}^{\prime}=\mathbf{s}+\mathbf{e}_{i}$. This completes the proof of surjectivity.
@@
This quantity is always at least $\operatorname{rk}_{P}(i)$ because $P$ is loopless; therefore, $\overline{(E_{i}\setminus e)\cup e^{\prime}}=\overline{E_{i}}$. The proof is now complete. ∎
@@
Define $\mathcal{L}_{P}:=\mathcal{L}_{P,\mathbf{n}}$, with $\mathbf{n}$ any cage for $P$. This definition does not depend on the choice of cage because we may assume that $P$ is loopless by 3.17, then repeatedly apply 3.16 to produce an isomorphism $\mathcal{L}_{P,\mathbf{n}}\cong\mathcal{L}_{P,(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))}$. We will henceforth assume that $\mathbf{n}=(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))$.
Statements i and iii are the content of 3.13; we now move to proving ii. To construct $P^{\mathrm{sim}}$: let $(P^{\prime},\mathbf{n}^{\prime})=R_{i}(P\setminus L,\pi_{E\setminus L}(\mathbf{n}))$, where $L$ is the set of loops of $P$ and $i\in E\setminus L$ is any element such that $\{i\}$ is not a flat. If the resulting polymatroid $P^{\prime}$ is simple, then take $P^{\mathrm{sim}}:=P^{\prime}$; otherwise, repeat these steps, replacing $(P,\mathbf{n})$ with $(P^{\prime},\mathbf{n}^{\prime})$. This process terminates because the cage shrinks at each step. If the cage is $\mathbf{0}$, then $P$ is the polymatroid of rank 0 on zero elements, which is simple; otherwise, the process stopped before the cage was equal to $\mathbf{0}$, meaning that a simple polymatroid was produced.
Lemmas 3.16 and 3.17 show that $\mathcal{L}_{P^{\mathrm{sim}}}\cong\mathcal{L}_{P}$, and 3.15 implies $P^{\mathrm{sim}}$ is the unique simple polymatroid with this property. This completes the proof of 1.1. ∎
@@
If $P$ is loopless and realized by a subspace arrangement $V_{1},\ldots,V_{N}\subset V$, then $P$ fails to be simple if and only if $V_{i}\subset V_{j}\subsetneq V$ for some $i\neq j$. In this case, a step of the “simplification process” used to construct $P^{\mathrm{sim}}$ in the proof of 1.1 corresponds to replacing $V_{i}$ with $V_{i}+\ell$, where $\ell\subset V$ is a generic line, then deleting loops. Alternatively, one may view this as replacing $V_{i}$ with its immediate successor $V_{i,\operatorname{codim}V_{i}-1}$ in a general flag containing $V_{i}$ (see 3.7 and 3.19).
@@
A poset $\mathcal{L}$ is the lattice of combinatorial flats of a polymatroid $P$ if and only if
@@
If $P$ is a polymatroid (which we may assume simple by 1.1ii), then the prescribed properties hold for $\mathcal{L}_{P}$ by 1.1i, 3.14, and 3.2.
@@
Plainly, the compositions $\varphi\circ\varphi^{\prime}$ and $\varphi^{\prime}\circ\varphi$ are increasing maps. To finish, we show that both $\varphi$ and $\varphi^{\prime}$ decrease rank. Let $\mathbf{s}$ be a combinatorial flat and let $F=\{i:s_{i}=n_{i}\}$, so that $\mathbf{s}^{\mathrm{geo}}=\sum_{i\in F}n_{i}\mathbf{e}_{i}$. Using submodularity of $\operatorname{rk}_{\mathcal{L}}$, our discussion of flats of $P$, and 3.2, we obtain
@@
4. Polymatroid Schubert varieties
We introduce and study the polymatroid Schubert variety $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ associated to a linear subspace $V\subset\mathbb{K}^{\mathbf{n}}$, with $\mathbb{K}$ an algebraically closed field of characteristic 0. Our construction is founded on work of Hassett and Tschinkel [HT99], which we review in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section 4.4, then in Section 4.5 compute the support of its class in the Chow ring of $\mathbb{P}^{\mathbf{n}}$. In Section 4.6, we discuss “polymatroid generality”, a condition that guarantees $Y_{V}$ consists of finitely many $V$-orbits. We assume this condition in Section 4.7, where we use the results of Section 4.5 to identify strata of $Y_{V}$, ending with a proof of 1.3. Finally, in Section 4.8, we give a presentation for the singular cohomology ring of $Y_{V}$ when $\mathbb{K}=\mathbb{C}$ and $V$ is polymatroid general.
@@ 4. Polymatroid Schubert varieties
[HT99, Proposition 3.7] Over an algebraically closed field of characteristic 0, there is a left action of $\mathbb{G}_{a}^{n}$ on $\mathbb{P}^{n}$ with finitely many orbits and such that the stabilizer of a general point of $\mathbb{P}^{n}$ is trivial. The action is unique up to isomorphism of $\mathbb{G}_{a}^{n}$-variety structures.
To make matters concrete, choose coordinates $a_{1},\ldots,a_{n}$ and $b_{0},\ldots,b_{n}$ on $\mathbb{G}_{a}^{n}$ and $\mathbb{P}^{n}$, respectively. The action on $\mathbb{P}^{n}$ is given by the faithful representation $\rho_{n}:\mathbb{G}_{a}^{n}\to\operatorname{Aut}(\mathbb{P}^{n})$,
@@
$B_{i}(x_{1},\ldots,x_{i})=\sum_{j_{1}+2j_{2}+\cdots+ij_{i}=i}\frac{x_{1}^{j_{1}}x_{2}^{j_{2}}\cdots x_{i}^{j_{i}}}{j_{1}!j_{2}!\cdots j_{i}!}\quad\text{for $i\geq 1$.}$
@@
$S_{k}\coloneqq\{a_{1}=a_{2}=\cdots=a_{k}=0\}\subset\mathbb{G}_{a}^{n},$
@@
Consider now the semidirect product $\mathbb{G}_{a}^{n}\rtimes\mathbb{G}_{m}$, in which $\mathbb{G}_{m}$ acts on $\mathbb{G}_{a}^{n}$ by $t\cdot\mathbf{a}=(ta_{1},t^{2}a_{2},\ldots,t^{n}a_{n})$. If we define
@@
in which the image of $\lambda_{n}$ normalizes the image of $\rho_{n}$. The orbits of $\mathbb{G}_{a}^{n}\rtimes\mathbb{G}_{m}$ on $\mathbb{P}^{n}$ are the same as those of $\mathbb{G}_{a}^{n}$, but each $\mathbb{G}_{a}^{n}$-orbit $O_{i}$ contains a unique $\mathbb{G}_{m}$-fixed point. In coordinates, the fixed point is $[\mathbf{e}_{n+1-i}]$. Consequently, we may canonically identify $O_{i}$ with $\mathbb{G}_{a}^{n}/S_{i}$ via the map $\mathbf{a}\mapsto\rho_{n}(\mathbf{a})\cdot[\mathbf{e}_{n+1-i}]$.
@@
We construct polymatroid Schubert varieties, bootstrapping from the Hassett-Tschinkel action of Section 4.1. The Hassett-Tschinkel maps $\rho_{n_{i}}:\mathbb{K}^{n_{i}}\to\operatorname{Aut}(\mathbb{P}^{n_{i}})$ assemble to a map $\rho:\mathbb{K}^{\mathbf{n}}\to\operatorname{Aut}(\mathbb{P}^{\mathbf{n}})$, allowing $\mathbb{K}^{\mathbf{n}}$ to act on $\mathbb{P}^{\mathbf{n}}$. For each multiset $\mathbf{s}\leq\mathbf{n}$, there is an $|\mathbf{s}|$-dimensional orbit $O_{\mathbf{s}}:=O_{s_{1}}\times\cdots\times O_{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$, whose closure is a product of projectivized coordinate subspaces $\mathbb{P}^{\mathbf{s}}:=\mathbb{P}^{s_{1}}\times\cdots\times\mathbb{P}^{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$. All points in $O_{\mathbf{s}}$ have the same stabilizer, denoted by $S_{\mathbf{s}}$. The explicit description in Section 4.1 shows that $S_{\mathbf{s}}$ is the coordinate subspace of $\mathbb{K}^{\mathbf{n}}$ cut out by

### v2

@@ 3. Combinatorial flats
Our results on Schubert varieties of subspace arrangements depend only on Sections 3.1 and 3.2, so geometrically-inclined readers may go directly to Section 4 after reading these.
@@
A combinatorial flat of $(P,\mathbf{n})$ is a multiset $\mathbf{s}\in\mathbb{Z}_{\geq 0}^{E}$ such that there exists a flat $F$ of $\widetilde{P}$ satisfying $\mathbf{s}=(|F\cap E_{1}|,\ldots,|F\cap E_{N}|)$. (Rephrased in the language of Section 1.5, $\mathbf{s}$ is a combinatorial flat if it is represented by a flat of $\widetilde{P}$.)
@@
Suppose that the polymatroid $P$ is realized by a subspace arrangement $V_{1},\ldots,~{}V_{N}~{}\subset~{}V$. Rephrasing 2.10 in the language of arrangements, if $\mathbf{n}$ is a cage for $P$, then $\widetilde{P}$ is realized by any arrangement $\{H_{ij}:1\leq i\leq N,1\leq j\leq n_{i}\}$ in which $H_{i1},\ldots,H_{in_{i}}$ are general hyperplanes containing $V_{i}$. The lattice of flats $\mathcal{L}_{\widetilde{P}}$ has an action by a product of symmetric groups $\mathfrak{S}$, and the quotient $\mathcal{L}_{\widetilde{P}}/\mathfrak{S}$ is isomorphic to $\mathcal{L}_{P}$ (see also Section 3.1).
@@
The if direction follows from 2.12 and 2.13. For the “only if”, suppose that $\pi_{E\setminus i}(\mathbf{s})$ is a combinatorial flat of $(P\setminus i,\pi_{E\setminus i}(\mathbf{n}))$. We will show that $\mathbf{s}$ must be a flat by proving $\mathbf{s}=\overline{\mathbf{s}}$. By 3.2 and the hypothesis that $\overline{s}_{i}<n_{i}$, it is always the case that $s_{i}=\overline{s}_{i}$. Hence, it suffices to show that $s_{j}=\overline{s}_{j}$ for $j\neq i$.
By 2.12 and 2.13, there is a combinatorial flat $\mathbf{s}^{\prime}$ of $(P,\mathbf{n})$ with $\pi_{E\setminus i}(\mathbf{s}^{\prime})=\pi_{E\setminus i}(\mathbf{s})$. If $s^{\prime}_{i}=n_{i}$, then $\mathbf{s}<\mathbf{s}^{\prime}$, so $\mathbf{s}\leq\overline{\mathbf{s}}<\mathbf{s}^{\prime}$. Entrywise, this implies $s_{j}\leq\overline{s}_{j}\leq s_{j}^{\prime}=s_{j}$, which gives the desired equality.
@@
By the matroid case and 2.15, a set $S$ is a flat of $\widetilde{T_{F}P}$ if and only if $S$ is a flat of $\widetilde{P}$ that either contains $\cup_{i\in F}E_{i}$ or satisfies $\overline{S\cup i}\not\supset\cup_{i\in F}E_{i}$ for all $i\in\widetilde{E}\setminus S$. In the former case, $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)-1$; in the latter $\operatorname{rk}_{\widetilde{T_{F}P}}(S)=\operatorname{rk}_{\widetilde{P}}(S)$. Taking $\mathbf{s}$ to be the multiset represented by $S$ completes the proof. ∎
In the remainder of Section 3, we work up to proving 1.1, then provide axioms for combinatorial flats. These results are not required to understand Schubert varieties of subspace arrangements; readers interested solely in geometry may proceed to Section 4 forthwith.
@@
$\leq\operatorname{rk}_{\widetilde{P}}(S)+\operatorname{rk}_{\widetilde{P}}(S^{\prime})=\operatorname{rk}_{P}(\mathbf{s})+\operatorname{rk}_{P}(\mathbf{s}^{\prime}).\qed$
@@
$\mathcal{L}_{P,\mathbf{n}}$ is a semimodular lattice, with join and meet given by $\vee$ and $\wedge$, respectively. Moreover, $\mathcal{L}_{P,\mathbf{n}}$ is top heavy and $|\mathcal{L}_{P,\mathbf{n}}^{k}|\leq|\mathcal{L}_{P,\mathbf{n}}^{k+1}|$ when $k\leq d/2$.
@@
where $S$ is any flat representing $\mathbf{s}$. By [BHM+20, Theorem 1.1], there is an injective linear $L:H^{k}(\widetilde{P})\to H^{d-k}(\widetilde{P})$ that commutes with the action of $\mathfrak{S}$. Consequently, the restriction of $L$ to $H^{k}(\widetilde{P})^{\mathfrak{S}}$ gives a linear injection
@@
and taking dimensions yields top-heaviness. Also by [BHM+20, Theorem 1.1], there are $\mathfrak{S}$-equivariant linear injections $H^{0}(\widetilde{P})\hookrightarrow H^{1}(\widetilde{P})\hookrightarrow\ldots\hookrightarrow H^{\lfloor d/2\rfloor}(\widetilde{P})$, which restrict to linear injections
$H^{k}(P,\mathbf{n})\cong H^{k}(\widetilde{P})^{\mathfrak{S}}\lhook\joinrel\xrightarrow{\;L\;}H^{k+1}(\widetilde{P})^{\mathfrak{S}}\cong H^{k+1}(P,\mathbf{n}),$
which proves the remaining statement. ∎
Generalizations of Lemmas 3.11 and 3.12 for matricubes appear in [AG24, Theorem 3.4] (see 3.8). With Lemmas 3.11 and 3.12 in hand, we have established 1.1 parts (i), (iii), and (iv). We now devote our attention to 1.1(ii), which states that $P$ is determined by $\mathcal{L}_{P}$ up to simplification.
An element of a lattice is join-irreducible if it cannot be written as a join of strictly smaller elements.
@@
By 3.13, each $i\in E$ corresponds to a maximal join-irreducible $x_{i}$ of $\mathcal{L}_{P,\mathbf{n}}$. The rank function of $P$ is recovered by $\operatorname{rk}_{P}(S)=\operatorname{rk}_{P,\mathbf{n}}(\vee_{i\in S}x_{i})$. ∎
@@
is an isomorphism. Multisets in the image of $\psi$ are combinatorial flats of $R_{i}(P,\mathbf{n})$ by Lemmas 2.17 and 2.12.
For surjectivity, suppose that $\mathbf{s}\in\mathcal{L}_{R_{i}(P,\mathbf{n})}$ is represented by a flat $S$ of $\widetilde{R_{i}(P,\mathbf{n})}$. We will show there is $\mathbf{s}^{\prime}\in\mathcal{L}_{P,\mathbf{n}}$ with $\psi(\mathbf{s}^{\prime})=\mathbf{s}$. By Lemmas 2.17 and 2.12, there is $e\in E_{i}\setminus S$ such that $S\cup e$ or $S$ is a flat of $\widetilde{P}$. If $S$ is a flat of $\widetilde{P}$, then $s_{i}<n_{i}$, so we may take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, suppose $S\cup e$ is a flat of $\widetilde{P}$. If $s_{i}+1<n_{i}$, then by 3.2, $\mathbf{s}$ is a combinatorial flat of $\widetilde{P}$, so we may again take $\mathbf{s}^{\prime}=\mathbf{s}$. Otherwise, if $S\cup e$ is a flat of $\widetilde{P}$ and $s_{i}+1=n_{i}$, then take $\mathbf{s}^{\prime}=\mathbf{s}+\mathbf{e}_{i}$. This completes the proof of surjectivity.
@@
This quantity is always at least $\operatorname{rk}_{P}(i)$ because $P$ is loopless; therefore, $\overline{(E_{i}\setminus e)\cup e^{\prime}}=\overline{E_{i}}$. ∎
@@
Define $\mathcal{L}_{P}:=\mathcal{L}_{P,\mathbf{n}}$, with $\mathbf{n}$ any cage for $P$. This definition does not depend on the choice of cage because we may assume that $P$ is loopless by 3.16, then repeatedly apply 3.15 to produce an isomorphism $\mathcal{L}_{P,\mathbf{n}}\cong\mathcal{L}_{P,(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))}$. We will henceforth assume that $\mathbf{n}=(\operatorname{rk}_{P}(1),\ldots,\operatorname{rk}_{P}(N))$.
Statements i and iv are the content of 3.12; we now move to proving ii. To construct $P^{\mathrm{sim}}$: let $(P^{\prime},\mathbf{n}^{\prime})=R_{i}(P\setminus L,\pi_{E\setminus L}(\mathbf{n}))$, where $L$ is the set of loops of $P$ and $i\in E\setminus L$ is any element such that $\{i\}$ is not a flat. If the resulting polymatroid $P^{\prime}$ is simple, then take $P^{\mathrm{sim}}:=P^{\prime}$; otherwise, repeat these steps, replacing $(P,\mathbf{n})$ with $(P^{\prime},\mathbf{n}^{\prime})$. This process terminates because the cage shrinks at each step. If the cage is $\mathbf{0}$, then $P$ is the polymatroid of rank 0 on zero elements, which is simple; otherwise, the process stopped before the cage was equal to $\mathbf{0}$, meaning that a simple polymatroid was produced.
Lemmas 3.15 and 3.16 show that $\mathcal{L}_{P^{\mathrm{sim}}}\cong\mathcal{L}_{P}$, and 3.14 implies $P^{\mathrm{sim}}$ is the unique simple polymatroid with this property. This completes the proof of 1.1. ∎
@@
If $P$ is loopless and realized by a subspace arrangement $V_{1},\ldots,V_{N}\subset V$, then $P$ fails to be simple if and only if $V_{i}\subset V_{j}\subsetneq V$ for some $i\neq j$. In this case, a step of the “simplification process” used to construct $P^{\mathrm{sim}}$ in the proof of 1.1 corresponds to replacing $V_{i}$ with $V_{i}+\ell$, where $\ell\subset V$ is a generic line, then deleting loops. Alternatively, one may view this as replacing $V_{i}$ with its immediate successor $V_{i,\operatorname{codim}V_{i}-1}$ in a general flag containing $V_{i}$ (see 3.7 and 3.18).
@@
A graded poset $\mathcal{L}$ is the lattice of combinatorial flats of a polymatroid $P$ if and only if
@@
If $P$ is a polymatroid (which we may assume simple by 1.1ii), then the prescribed properties hold for $\mathcal{L}_{P}$ by 1.1i, 3.13, and 3.2.
@@
Plainly, the compositions $\varphi\circ\varphi^{\prime}$ and $\varphi^{\prime}\circ\varphi$ are increasing maps (i.e. $(\varphi\circ\varphi^{\prime})(x)\geq x$ and $(\varphi^{\prime}\circ\varphi)(\mathbf{s})\geq\mathbf{s}$). To finish, we show that both $\varphi$ and $\varphi^{\prime}$ decrease rank. Let $\mathbf{s}$ be a combinatorial flat and let $F=\{i:s_{i}=n_{i}\}$, so that $\mathbf{s}^{\mathrm{geo}}=\sum_{i\in F}n_{i}\mathbf{e}_{i}$. Using submodularity of $\operatorname{rk}_{\mathcal{L}}$, our discussion of flats of $P$, and 3.2, we obtain
@@
4. Schubert varieties of subspace arrangements
We introduce and study the Schubert variety $Y_{V}\subset\mathbb{P}^{\mathbf{n}}$ associated to a linear subspace $V\subset\mathbb{K}^{\mathbf{n}}$, with $\mathbb{K}$ an algebraically closed field of characteristic 0. Our construction is founded on work of Hassett and Tschinkel [HT99], which we review in Sections 4.1 and 4.2. We construct $Y_{V}$ in Section 4.4, then in Section 4.5 compute the support of its class in the Chow ring of $\mathbb{P}^{\mathbf{n}}$. In Section 4.6, we discuss “polymatroid generality”, a condition that guarantees $Y_{V}$ consists of finitely many $V$-orbits. We assume this condition in Section 4.7, where we use the results of Section 4.5 to identify strata of $Y_{V}$, ending with a proof of 1.3. Finally, in Section 4.8, we give a presentation for the singular cohomology ring of $Y_{V}$ when $\mathbb{K}=\mathbb{C}$ and $V$ is polymatroid general.
@@ 4. Schubert varieties of subspace arrangements
[HT99, Proposition 3.7] Over an algebraically closed field of characteristic 0, there is a left action of $\mathbb{G}_{a}^{n}$ on $\mathbb{P}^{n}$ with finitely many orbits.The action is unique up to isomorphism of $\mathbb{G}_{a}^{n}$-variety structures.
To make matters concrete, choose coordinates $a_{1},\ldots,a_{n}$ and $b_{0},\ldots,b_{n}$ on $\mathbb{G}_{a}^{n}$ and $\mathbb{P}^{n}$, respectively. The action on $\mathbb{P}^{n}$ is given by the faithful representation $\rho_{n}:\mathbb{G}_{a}^{n}\to\operatorname{Aut}(\mathbb{P}^{n})=\operatorname{PGL}_{n+1}$,
@@
$B_{i}(x_{1},\ldots,x_{i})=i!\sum_{j_{1}+2j_{2}+\cdots+ij_{i}=i}\frac{1}{j_{1}!j_{2}!\cdots j_{i}!}\left(\frac{x_{1}}{1!}\right)^{j_{1}}\left(\frac{x_{2}}{2!}\right)^{j_{2}}\cdots\left(\frac{x_{i}}{i!}\right)^{j_{i}}\quad\text{for $i\geq 1$.}$
@@
$\operatorname{St}_{k}\coloneqq\{a_{1}=a_{2}=\cdots=a_{k}=0\}\subset\mathbb{G}_{a}^{n},$
@@
Consider now the semidirect product $\mathbb{G}_{a}^{n}\rtimes\mathbb{G}_{m}$ (with $n\geq 1$), in which $\mathbb{G}_{m}$ acts on $\mathbb{G}_{a}^{n}$ by $t\cdot\mathbf{a}=(ta_{1},t^{2}a_{2},\ldots,t^{n}a_{n})$. If we define
@@
in which the image of $\lambda_{n}$ normalizes the image of $\rho_{n}$. The orbits of $\mathbb{G}_{a}^{n}\rtimes\mathbb{G}_{m}$ on $\mathbb{P}^{n}$ are the same as those of $\mathbb{G}_{a}^{n}$, but each $\mathbb{G}_{a}^{n}$-orbit $O_{i}$ contains a unique $\mathbb{G}_{m}$-fixed point. In coordinates, the fixed point is $[\mathbf{e}_{n+1-i}]$. Consequently, we may canonically identify $O_{i}$ with $\mathbb{G}_{a}^{n}/\operatorname{St}_{i}$ via the map $\mathbf{a}\mapsto\rho_{n}(\mathbf{a})\cdot[\mathbf{e}_{n+1-i}]$.
@@
We construct Schubert varieties of subspace arrangements, bootstrapping from the Hassett-Tschinkel action of Section 4.1. Identifying $\mathbb{G}_{a}^{n_{i}}$ and $\mathbb{K}^{n_{i}}$, the Hassett-Tschinkel maps $\rho_{n_{i}}:\mathbb{K}^{n_{i}}\to\operatorname{Aut}(\mathbb{P}^{n_{i}})$ assemble to a map $\rho:\mathbb{K}^{\mathbf{n}}\to\operatorname{Aut}(\mathbb{P}^{\mathbf{n}})$, allowing $\mathbb{K}^{\mathbf{n}}$ to act on $\mathbb{P}^{\mathbf{n}}$. For each multiset $\mathbf{s}\leq\mathbf{n}$, there is an $|\mathbf{s}|$-dimensional orbit $O_{\mathbf{s}}:=O_{s_{1}}\times\cdots\times O_{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$, whose closure is a product of projectivized coordinate subspaces $\mathbb{P}^{\mathbf{s}}:=\mathbb{P}^{s_{1}}\times\cdots\times\mathbb{P}^{s_{N}}\subset\mathbb{P}^{\mathbf{n}}$. All points in $O_{\mathbf{s}}$ have the same stabilizer, denoted by $\operatorname{St}_{\mathbf{s}}$. The explicit description in Section 4.1 shows that $\operatorname{St}_{\mathbf{s}}$ is the coordinate subspace of $\mathbb{K}^{\mathbf{n}}$ cut out by